OpenStudy (aaliyahfairgood):

Johnny Mac's Sporting Goods bought a baseball glove from Rawlings Sporting Goods for $66.00. They want to mark up the glove 70% on selling price. What should Johnny's sell the glove for? A. $179.82 B. $113.22 C. $246.42 D. $220.20

OpenStudy (aaliyahfairgood):

>>>mark-up price + original price >>66(70%) + original price = >46.2 +66.0 = >112.2

OpenStudy (aaliyahfairgood):

Am I doing this wrong?

OpenStudy (texaschic101):

66 + 0.70(66) = 112.20 you got what I got...hmm...not an answer choice

OpenStudy (aaliyahfairgood):

I hate questions like this -_-

OpenStudy (texaschic101):

@Directrix ....what am I missing here

OpenStudy (aaliyahfairgood):

idk lol I just wanted to know if I am missing something too

OpenStudy (texaschic101):

is this a percent increase problem ?

Directrix (directrix):

66.00 + 70% (66) for a 70% increase although that was not specified.

OpenStudy (aaliyahfairgood):

Business and Finance

Directrix (directrix):

Markup is the difference between the cost of a good or service and its selling price. A markup is added onto the total cost incurred by the producer of a good or service in order to create a profit. The total cost reflects the total amount of both fixed and variable expenses to produce and distribute a product.

Directrix (directrix):

That sounds like this to me: 66.00 + 70% (66)

Directrix (directrix):

$ 112. 20

OpenStudy (aaliyahfairgood):

so the answer choices are wrong then ? 0.o

OpenStudy (aaliyahfairgood):

that could make it B. 113.22 or D. 222.20

OpenStudy (aaliyahfairgood):

Thank you guys, both of you

OpenStudy (texaschic101):

(113.22 - 66) / 66 = 47.22/66 = 0.715 = 71.5%......not 70%....I do not know...I could be looking at this problem the wrong way

OpenStudy (texaschic101):

thats it directrix...

OpenStudy (aaliyahfairgood):

220.20

OpenStudy (texaschic101):

I would go with 220.20

OpenStudy (texaschic101):

although I am not sure where the extra 20 cents comes in

OpenStudy (aaliyahfairgood):

Thank You ^.^